### v1

# Luminous Fast Blue Optical Transients as very massive star core-collapse events

Abstract
Context. Luminous Fast Blue Optical Transients (LFBOTs) are rare extragalactic events of unknown origin. Tidal disruptions of white dwarfs by intermediate mass black holes, mergers of black holes and Wolf-Rayet stars, and failed supernovae are among the suggested explanations.
Aims. In this paper, we explore the viability of very massive star core-collapse events as the origin of LFBOTs. The appeal of such a model is that the formation of massive black holes via core collapse may yield observational signatures that can match the disparate lines of evidence that point towards both core-collapse and tidal disruption origins for LFBOTs.
Methods. We explore the formation rate of massive black holes in binary population synthesis models, and compare the metallicities of their progenitors with the observed metallicities of LFBOT host galaxies. We further examine the composition, mass loss rates and fallback masses of these stars, placing them in the context of LFBOT observations.
Results. The formation rate of black holes with masses greater than 30–40 M⊙ is found to be similar to the observed LFBOT rate. The stars producing these black holes are biased to low metallicity (Z0.3 Z⊙), are H and He-poor and have dense circumstellar media. However, some LFBOTs have host galaxies with higher metallicities than predicted, and others have denser local environments (plausibly due to late stage mass loss not captured in the models). We find that long-lived emission from an accretion disc (as implicated in the prototypical LFBOT AT 2018cow) can plausibly be produced in these events.
Conclusions. We conclude that (very) massive star core-collapse, based on the above criteria, is a plausible explanation for LFBOTs. The preferred progenitors for LFBOTs in this scenario overlap with those predicted to produce super-kilonovae. We therefore suggest that LFBOTs are promising targets to search for super-kilonovae, and that they may contribute non-negligibly to the r-process enrichment of galaxies.
Key Words.:
Supernovae: general – Stars: black holes – Stars: massive1 Introduction
Luminous Fast Blue Optical Transients (LFBOTs), or Cow-like transients after the prototypical AT 2018cow (Prentice et al., 2018; Perley et al., 2019), are rapidly evolving transients with peak optical-ultraviolet luminosities rivalling superluminous supernovae (SLSNe). The early optical-UV spectra are well described by a hot, featureless black body (T20000 K), lacking H and He lines initially. They have decay rates of 0.3 magnitudes/day, ruling out a significant contribution from 56Ni-powered emission. They are also accompanied by bright X-ray (and at late times, radio) emission. The late-time radio emission is attributed to self-absorbed synchrotron emission from an expanding blastwave in a dense circumstellar medium (CSM).
A modest sample of LFBOTs has now been identified in addition to AT 2018cow: ZTF18abvkwla (AT2018lug, Ho et al., 2020), CSS161010 (Coppejans et al., 2020; Gutiérrez et al., 2024), ZTF20acigmel (AT2020xnd, Perley et al., 2021; Bright et al., 2022), AT2020mrf (Yao et al., 2022), AT2022tsd (Matthews et al., 2023), AT2023fhn (Chrimes et al., 2024b, a) and AT2024wpp (Pursiainen et al., 2025; LeBaron et al., 2025; Nayana et al., 2025), as well as the slower-evolving and radio-faint, but otherwise similar, AT2024puz (Somalwar et al., 2025) and several other candidates which lack extensive multi-wavelength follow-up. Some transients display very blue colours, evolve rapidly at early times and produce bright X-ray emission, similar to LFBOTs, but are followed by broad lined type Ic SNe (e.g. the Einstein Probe-discovered transients EP2404014A and EP250108A, van Dalen et al., 2025; Eyles-Ferris et al., 2025; Rastinejad et al., 2025). However, these events likely do not share the same progenitors as LFBOTs, given that they successfully launched supernovae, and we henceforth focus solely on the SN-less, Cow-like LFBOTs, whose origins remain unknown. The volumetric rate of LFBOTs is estimated to be 0.1% of the core-collapse SN rate (Ho et al., 2023b).
Two categories of model to explain LFBOTs have emerged as most likely (e.g. Perley et al., 2019). The first invokes tidal disruption events (TDEs) of H and He poor, compact stars (e.g. a white dwarfs) by intermediate mass black holes (IMBHs, Kuin et al., 2019). This model is challenged primarily due to the dense CSM observed in LFBOTs. Although such a environment is possible in TDEs involving stars disrupted by supermassive black holes, it is less clear whether white dwarf plus IMBH TDEs could produce similar circumstellar density profiles (Margutti et al., 2019; Linial & Quataert, 2024). The second category of proposals are broadly themed around stellar mass black hole accretion, either in mergers/collisions (Metzger, 2022; Grichener, 2025; Tsuna & Lu, 2025) or core-collapse events (e.g. Kashiyama & Quataert, 2015; Margutti et al., 2019; Gottlieb et al., 2022).
Sufficiently massive stars are expected to directly form black holes without launching a supernova. The landscape of explodability is thought to be complex, with neutron star and some black hole-forming events capable of launching successful supernovae, and other black hole-forming core-collapses resulting in direct black hole formation without any significant ejecta (e.g. Heger et al., 2003; Ertl et al., 2020). Observational evidence for ‘quiet’ black hole formation comes from a lack of massive stars M⊙ in pre-supernovae imaging, and the disappearance of supergiant stars without evidence for a supernova (Smartt et al. 2009; Williams et al. 2014; Reynolds et al. 2015; Adams et al. 2017; Beasor et al. 2024). Some LFBOT models invoke such a scenario, where a massive star collapses to a black hole without a successful supernova (a ‘failed supernova’), followed by accretion of the remaining (‘leftover’) material onto the nascent BH (e.g. Fernández et al., 2018; Quataert et al., 2019). The failed supernova hypothesis resembles early collapsar gamma-ray burst (GRB) models (e.g. Woosley, 1993; MacFadyen & Woosley, 1999; Fryer et al., 1999), which also feature accreting black holes born in core-collapse, but without a successful (or a weak) supernova. In LFBOTs there is no evidence for a supernova, but unlike in collapsar GRBs, no gamma-ray emission has been observed, and the outflows are only mildly relativistic (e.g. Coppejans et al., 2020).
While some fast evolving optical transients can be explained solely with circumstellar interactions (Pellegrino et al., 2022), in LFBOTs there is evidence for central engine activity. AT 2018cow showed short timescale variability in its light-curve (Ho et al., 2019; Margutti et al., 2019; Pasham et al., 2021), as well as UV and X-ray emission lasting at least several years post-explosion (Sun et al., 2022, 2023; Inkenhaag et al., 2023; Chen et al., 2023; Migliori et al., 2024). Magnetar central engines have been proposed (Fang et al., 2019; Mohan et al., 2020; Liu et al., 2022; Li et al., 2024), but struggle to explain both the early-time emission and late-time optical/UV plateau (Chen et al., 2023). Assuming instead that LFBOTs are black hole (BH) powered, estimates for the BH mass in such a scenario have been derived from quasi period oscillations (Pasham et al., 2021) and accretion disc modelling (Inkenhaag et al., 2023; Migliori et al., 2024; Cao et al., 2024) yielding masses which span the upper end of the stellar mass BH distribution into the intermediate BH mass regime. The observation of late-time, short duration giant optical flares from AT 2022tsd (Ho et al., 2023a) may also be consistent with a highly variable accretion rate, either around the black hole formed in the event or a companion black hole (Lazzati et al., 2024). Similar flaring behaviour has not been found in AT20224wpp (Ofek et al., 2025).
Given the high BH mass estimates in LFBOTs, and their discovery thus far in non-nuclear regions of low-metallicity, star-forming galaxies, the primary aim of this paper is determine whether massive BH formation in failed supernovae is consistent with the volumetric rate and environments of LFBOTs. In Section 2 we compare population synthesis predictions for the rate of BH formation as a function of mass with the observed LFBOT rate. We then investigate in Section 3 whether the metallicity bias associated with the formation of the most massive stellar mass BHs is consistent with LFBOT environments. We explore in Section 4 whether the selected stellar models can plausibly reproduce the key characteristics of LFBOT emission, with a discussion and conclusions following in Sections 5 and 6.
2 Black hole formation and LFBOT event rates
We first ask the question: what is the volumetric formation rate of stellar-mass, core-collapse black holes with mass MBH,min, and which value of MBH,min best reproduces the LFBOT rate? To explore this, we couple population synthesis with a model for the metallicity-dependent cosmic star formation rate history.
For the population synthesis we use the Binary Population and Spectral Synthesis (bpass) models (v2.2.1, Eldridge et al., 2017; Stanway & Eldridge, 2018). These are a publicly available grid of pre-calculated binary stellar evolution models, in which the primary evolution is followed in detail, while the secondary is evolved with the rapid evolution prescriptions of Hurley et al. (2002). Models are provided at 13 metallicities and at each metallicity, each stellar model is weighted according to observed binary parameter distributions (Moe & Di Stefano, 2017). We adopt the model set with a Kroupa (2001) broken power-law initial mass function (IMF), with a minimum mass of 0.1 M⊙ and a maximum of 300 M⊙, a break mass of 0.5 M⊙ and a slope of () below (above) the break. At each metallicity , the weighting of each model corresponds to the number of systems resembling that model expected in a stellar population of M⊙. For full details about the binary stellar models and the code, we refer the reader to the bpass v2 release paper Eldridge et al. (2017).
For the metallicity-dependent cosmic star formation rate history, CSFH, we adopt the model of Langer & Norman (2006). We seed the population synthesis models according to this CSFH such that the volumetric birth rate of each system (bpass model) at each redshift and metallicity is proportional to the product of the star formation rate density SFRD and the model weighting (defined by the IMF and binary parameter distributions as described above).
We adopt the standard bpass assumptions that core-collapse occurs when, at the end of the model, the total mass exceeds 1.5 M⊙, the CO core mass exceeds 1.38 M⊙ and the ONe core mass is non-zero. Remnant masses are a pre-calculated output of the models. They are determined by injecting 1051 erg of energy, calculating the mass in the outer layers which can be lifted to infinity by this energy injection, and taking the remaining mass as the remnant mass (for full details see Eldridge & Tout, 2004; Eldridge et al., 2017). We classify all core-collapse remnants with M3 M⊙ as neutron stars, and those heavier as black holes. We assume that all NS-forming events produce successful supernovae, and that BH-forming events do not. Pair-instability supernovae (PISNe) are deemed to occur if the final carbon-oxygen core mass exceeds 60 M⊙ and the final helium core mass is 133 M⊙. In this case no remnant is left behind (Heger & Woosley, 2002). Pulsational pair instability supernovae (PPISNe) are accounted for by adopting the remnant mass prescription of Farmer et al. (2019) for final carbon-oxygen masses between 38 and 60 M⊙, following Briel et al. (2023). The true landscape of explodability and remnant formation is more complex, but to first order our approximations are expected to be reasonable (e.g. Smartt et al., 2009; Smartt, 2009; Fryer et al., 2012; Ertl et al., 2020; Sukhbold & Adams, 2020; Patton & Sukhbold, 2020; Patton et al., 2022; Laplace et al., 2025; Ugolini et al., 2025; Steinwandel & Goldberg, 2025; The LIGO Scientific Collaboration et al., 2025).
The remnant formation time for each stellar model is the seed time, plus the age of the model at core-collapse (although for these short-lived stars, this is negligible compared with cosmological timescales). In this way, we construct a metallicity-weighted supernova rate and BH formation rate as a function of redshift. Since all (confirmed) LFBOTs thus far have redshifts , we compare the mean rate of supernovae and BH-forming events over this redshift range, for different values of MBH,min. Specifically, we calculate the mean ratio of the supernova rate and the formation rate of BHs with mass MBH,min or greater. As the form of CSFH is uncertain, we repeat the process with two extreme variations of the CSFH. These are defined by shifting the Langer & Norman (2006) model by 0.2 dex in 12log(O/H), which captures the approximate envelope of possible CSFHs found by Chruslinska & Nelemans (2019).
The results are shown in Figure 1. Using the fiducial CSFH we find that the birth rate of BHs at with masses 38 M⊙ is 0.1% of the CCSN rate, matching LFBOTs. For a CSFH of the same form but peaking at a metallicity 0.2 dex lower in 12log(O/H), we find a higher BH mass threshold of 41 M⊙ is required, because there are more high-mass BHs being formed overall with respect to CCSNe. The reverse is true for a CSFH which peaks 0.2 dex higher, yielding a threshold of 33 M⊙. Therefore, varying the CSFH changes the absolute number of BHs being born (although not the general shape of the BH mass distribution, van Son et al., 2023), and the number formed with respect to CCSNe.
Above mass cuts of 45 M⊙, the ratio of birth rates remains constant, because the cut has moved into the PISN mass gap, and only black holes with masses above the gap are contributing. In summary, we find that the birth rate of BHs heavier than several tens of Solar masses is similar to the LFBOT rate, and that this result is robust against uncertainties in the CSFH.
3 Metallicity dependence and LFBOT host galaxies
| LFBOT | Z/Z⊙ | Reference |
| AT 2018cow | 0.470.01 | Lyman et al. (2020)⋆ |
| ZTF18abvkwla | 0.37 | Ho et al. (2020) |
| CSS161010 | 0.15 | Gutiérrez et al. (2024) |
| ZTF20acigmel | 0.15 | Perley et al. (2021)† |
| AT2020mrf | 0.27 | Yao et al. (2022) |
| AT2023fhn | 0.080.02 | Chrimes et al. (2024a) |
| AT2024wpp | 0.53 | Pursiainen et al. (2025)† |
⋆ - IFU metallicity measurement, resolves the host giving a local (few hundred parsec scale) value. † - Metallicity derived from the host’s total stellar mass using equation 3 of Tremonti et al. (2004). Uncertainties are solely propagated from uncertainties on the galaxy stellar mass, and do not include scatter in the relation.
Given a black hole mass cut MBH,min which yields a BH formation rate matching the observed LFBOT rate, for an assumed CSFH, we can next ask: what is the (CSFH weighted) metallicity distribution of the progenitors of these black holes, and how does this compare with the host galaxies of LFBOTs? We collate the available information in the literature on LFBOT host metallicities in Table 1. Four of the seven have host-averaged metallicities derived from spectral energy distribution (SED) fitting. In one case (AT 2018cow) there is integral field unit data available, enabling a more local (resolution of a few hundred parsec, Lyman et al., 2020) measurement of the metallicity. For the remaining two, no direct measurement has been published, so we obtain an approximate value through the galaxy mass-metallicity relation (Tremonti et al., 2004). We adopt Z by mass fraction for Solar metallicity, but note that this specific choice has no impact on the comparisons or analysis since both the observed and synthetic metallicities are scaled by the same factor.
In Figure 2 we compare the number of black holes more massive than MBH,min born per M⊙ of star formation, weighted by the mean metallicity distribution for each of the three CSFHs adopted, with the measured distribution of LFBOT host or environmental metallicities. There is a strong preference for BHs with masses greater than 38 M⊙ to be born in stellar populations with metallicity less than 0.3 Z⊙. Below this threshold, the distributions are relatively constant: this reflects that the fact that (i) we are tailoring our mass cut to yield a BH formation rate that is 0.1% of the CCSN rate for each CSFH, and (ii) that the CCSN rate itself is not substantially varying with the CSFH variations. This is because NS formation efficiency, unlike BH formation efficiency, does not sharply decline at higher metallicities (e.g. van Son et al., 2025).
It is immediately apparent that LFBOT environments sample the ZZ⊙ region. The formation of BHs with MMBH,min across this range instead samples metallicities ZZ⊙. A KS-test between the LFBOT host values and the fiducial M⊙ distribution returns , formally rejecting the null hypothesis that the observed host values are consistent with the synthetic distribution. However, we note that the host metallicities are far from uniformly determined - with methods ranging from resolved IFU spectroscopy, probing the local environment within the host, to host-integrated metallicites to the mass-metallicity relation. This likely introduces errors/scatter which has not been accounted for, and host-integrated measurements will naturally be higher than the lowest metallicity environments within each host. There may also be pockets of low metallicity gas within and around star-forming galaxies, driving low-Z star-formation (e.g. Michałowski et al., 2015; Metha & Trenti, 2023). Nevertheless, the current information suggesting that LFBOTs can occur in environments with metallicities above the strong cut-off predicted is a challenge to the model. This is most notable for AT2018cow, whose metallicity measurement is integrated across the immediate (few hundred pc) environment (Lyman et al., 2020), but even in this case variations on smaller scales are possible.
4 Expectations for the emission
4.1 H and He-poor spectra
LFBOTs display H and He poor spectra at early times (e.g. Prentice et al., 2018). The progenitor star, or star being tidally disrupted, therefore must have a He or He-poor composition. In the TDE scenario, a white dwarf is favoured (Kuin et al., 2019). For a BH-stellar merger, the stellar object must be a Wolf-Rayet star (Metzger, 2022). In the scenario explored in this paper, the median H and He mass fractions in the ‘leftover mass’111I.e. the stellar mass leftover after remnant formation. In a successful supernova with no fallback, this would be the ejecta mass. of the models selected as progenitors of BHs with M⊙ are X0.003 and Y0.08 respectively (where the uncertainties are defined by the 16th and 84th percentiles). The distributions of Xsurf and Ysurf are shown in Figure 3. 69% of our selected models are classified as Wolf-Rayet stars given the Xsurf values and high surface temperatures of log10(T/K)4.45 (Eldridge et al., 2017). The outer layers of these stars therefore meet the criteria for being H and He poor, and successfully exploding stars with such low X and Y surface mass fractions are expected to produce type Ic supernovae (i.e. SNe without detectable H and He lines, e.g. Dessart et al., 2012; Eldridge et al., 2013; Chrimes et al., 2020).
4.2 Dense circumstellar media and radio emission
A key characteristic of LFBOTs is their slow-rising, luminous radio emission, attributed to self-absorbed synchrotron emission from a blast-wave expanding through the circumstellar medium (CSM). These blast-waves are mildly relativistic (e.g. Coppejans et al., 2020), and probe the CSM at radii of 1016–1017 cm. LFBOT radio emission is well described by blast-wave propagation through dense CSM profiles defined by values of 0.1-10, in units of M⊙ yr-1/1000 km s-1. For Wolf-Rayet wind speeds of order 1000 km s-1, radii of 1016–1017 cm corresponds to mass loss in the final decades before explosion. We calculate the mean final values for the selected models, noting that bpass only runs until the end of core carbon burning. Mass loss rates are provided as a bpass output. We adopt standard bpass wind speed prescriptions, namely those of Vink et al. (2001), and Nugis & Lamers (2000) for Wolf-Rayets (we refer the reader to Eldridge et al., 2017; Chrimes et al., 2022, for more details). We find a median final = (0.04) M⊙ yr-1/1000 km s-1. The full distribution of is shown in Figure 3. These mass-loss rate and wind speed values are consistent with the lower end of the values inferred from LFBOT radio observations, which span 0.1–10 (e.g. Yao et al., 2022). We note that mass loss in the final decades before core-collapse may be significant and eruptive, plausibly explaining the non- profiles observed (e.g. Smith, 2014). However, because the stellar evolution models used end after core carbon burning, a full modelling of the CSM around these stars on the relevant scales is beyond the scope of this paper.
4.3 Long-lived ultraviolet plateaus
Finally, we ask whether the selected stellar models can produce a long-lived tail of UV emission lasting at least several years after the event, as observed in AT2108cow (Sun et al., 2022, 2023; Chen et al., 2023; Inkenhaag et al., 2023). This has been interpreted as due to a long-lived accretion disc, reminiscent of the emission seen in TDEs (Inkenhaag et al., 2023; Mummery & van Velzen, 2025). To see whether our models can produce AT 2018cow-like UV plateaus through this mechanism, we first examine the remaining mass in the selected bpass models which survives outside the nascent BHs (i.e., the pre-core-collapse mass minus the initial BH mass). In Figure 4 we show 2D histograms of these masses at each metallicity, for the fiducial Langer & Norman (2006) CSFH. For more details on the BH mass distribution in bpass, we refer the reader to Eldridge et al. (2017), Ghodla et al. (2022) and Briel et al. (2023).
As outlined in Section 2, bpass calculates BH masses by injecting a standard supernova energy of 1051 erg to determine the mass of the material which is unbound. The remainder is the remnant mass. In a failed supernova, the ‘leftover’ material is not successfully ejected. It may therefore be reasonable to assume that all of the leftover material falls back and is accreted instead, in line with other prescriptions for final BH masses in the massive to very massive star regime, which have close to a one-to-one relation between the pre-core-collapse mass and final remnant mass (after fallback) for stripped envelope progenitors (Fryer et al., 2012; Briel et al., 2023).
With the BH mass, leftover mass, and angular momentum of the leftover mass at the radius of the pre-collapse stellar surface as inputs, we use the model of Mummery & Balbus (2020); Mummery et al. (2024) to calculate UV light-curves for every selected stellar model. The results for the F225W Wide Field Camera 3 (WFC3) Hubble Space Telescope UV filter are shown in Figure 6. We adopt two example scenarios. In the first scenario, all of the leftover material is accreted (as motivated above). This is the ‘100%’ case. We also explore a contrasting scenario in which only 10% of the leftover material is accreted (i.e. 90% is ejected, the 10% case). Massive stars in low metallicity environments such as the Magellanic Clouds are observed to be rotating at 10–20% of their critical velocity (e.g. Ramírez-Agudelo et al., 2013, 2015). In both scenarios, since stellar rotation is not explicitly tracked in bpass, we therefore assume that the star is rotating at 10% of the critical (break-up) velocity. The resulting light-curves primarily lie below the observed late-time F225W points for AT 2018cow, but with a substantial spread, and many models can match the luminosity of AT 2018cow (see also Omand et al. in prep).
We acknowledge that the failed SN interpretation of LFBOTs is predicated on the notion that some non-negligible fraction of the total stellar mass does not participate in the prompt BH formation and instead falls back onto the nascent BH. If the core-collapse of such massive stars actually results in the entire stellar mass being promptly engulfed by the event horizon, there will no material to accrete, ruling out failed supernovae as a viable channel for LFBOTs. However, in our fiducial models this is not the case (see Figure 4).
5 Discussion
5.1 LFBOT environments
In the failed SN scenario, we would expect LFBOTs to be strongly associated with star-forming regions, as the short-lived massive progenitors are not expected to travel far from their birth sites. Although the LFBOT sample is small, it appears as though LFBOTs occupy diverse environments within their hosts, with at least two events at moderate to large host-normalised offsets (Chrimes et al., 2024b; Pursiainen et al., 2025). The apparent preference for LFBOTs to occur in low metallicity environments may help in this regard: in general, galactic outskirts have lower metallicity than their inner regions (e.g. Pilkington et al., 2012; Ju et al., 2025). However, a similar argument could be made for e.g. core-collapse gamma-ray bursts, which are typically not observed at large offsets. Regardless of offsets, the host metallicities of LFBOTs extend above the sharp cut-off predicted for very massive star core-collapse events (see Sec 3). Furthermore, metallicities from SED fitting should be treated with caution (e.g. Nersesian et al., 2025). For example, the mass-metallicity relation for the host of AT 2023fhn suggests a metallicity of 0.5 Z⊙, versus the fitted value of 0.1 Z⊙. In any case, the metallicity distributions as presented in Fig. 2 are inconsistent, but the possibility of low-Z pockets on spatial scales smaller than probed by current measurements has the potential to resolve this discrepancy.
5.2 Initial mass function
We have adopted model weightings with an initial mass function (IMF) slope above 0.5 M⊙ of -2.35. Two other high-end slopes are available with the bpass outputs: -2.00 and -2.70. If the -2.70 slope is adopted, fewer high-mass black holes are born, and the cut-off MBH,min drops to 30 M⊙ such that the BH formation rate remains consistent with the LFBOT rate (0.1% of the CCSN rate). The number of neutron star progenitors also decreases, such that the SN rate decreases, reducing the amount by which MBH,min has to decrease in order to maintain 0.1% of the CCSN rate. If -2.00 is adopted, there are more black holes and the cut-off moves up to 40 M⊙, around the lower edge of the PISN mass gap. This range (i.e. 30–40 M⊙) represents the uncertainty in the minimum black hole mass required to match the LFBOT rate which arises from the choice of IMF, and is slightly larger than the uncertainty arising solely from the choice of CSFH. Nevertheless, the broad conclusion is that the formation rate of black holes with masses greater than 30–40 M⊙ is consistent with the LFBOT rate.
5.3 A potential source of r-process nucleosynthesis
Massive black holes and their accretion discs, such as those discussed here, have also been predicted as possible sites of ‘super-kilonovae’. In this scenario, the density in a high accretion disc becomes high enough to trigger r-process nucleosynthesis (Siegel et al., 2019, 2022; Dean & Fernández, 2024; Rastinejad et al., 2024; Agarwal et al., 2025). The possibility of such a channel has implications for where and when r-process elements are produced in the Universe (Nugent et al., 2025). Given the black hole and accreted masses involved in the failed supernova interpretation of LFBOTs, they may be promising sites for r-process production. The volumetric rate of LFBOTs in the redshift range considered in this work (Ho et al., 2023b, 0.1% of the CCSN rate, or 70 yr-1 Gpc-3) is compatible with the rate of super-kilonovae predicted by Siegel et al. (2022) of 10–100 yr-1 Gpc-3. We have shown that the failed-supernova interpretation of LFBOTs naturally favours events forming black holes greater than M⊙, while super-KNe have been predicted in events producing black holes greater than M⊙ (Siegel et al., 2022). If the true rate of LFBOTs is just a factor of few lower than the current 70 yr-1 Gpc-3 best estimate, then the BH mass cut required to match the LFBOT rate moves up to the pair instability mass gap (see Figure 1), and we have a 1:1 mapping between LFBOT and super-KN progenitors. In any case, there is surely substantial overlap in the populations. Furthermore, a near-infrared excess - a hallmark of radiative reprocessing by r-process elements - was observed on a timescale of weeks following both AT 2018cow (Perley et al., 2019) and AT2024wpp (Pursiainen et al., 2025). This has been described either as a dust echo or due to free-free opacity effects outside the optical photosphere in both events (Metzger & Perley, 2023; LeBaron et al., 2025), but no spectroscopic characterisation of this feature in an LFBOT has yet been obtained. Future (James Webb Space Telescope) near-infrared spectroscopic observations can distinguish between an r-process element-forming super-kilonova and a featureless dust echo or free-free emission.
5.4 Further implications
There are several other observations associated with LFBOTs which any model invoked to explain them must satisfy, assuming that LFBOTs are homogenous class. For example, extremely high velocity (), blue-shifted hydrogen lines were observed in the CSS161010 (Gutiérrez et al., 2024), perhaps most easily explained as out-flowing streams of material in a tidal disruption event, although these have only been observed in CSS161010 thus far. Polarimetric observations have found that LFBOTs start highly polarised and trend towards low polarisation on a timescale of days (Maund et al., 2023; Pursiainen et al., 2025). This can plausibly be explained in a failed supernova context by (i) an accretion disc producing the high early-time polarisation and (ii) a rapid transition to an spherical outflow (Pursiainen et al., 2025), perhaps caused by outflows from highly super-Eddington accretion (e.g. Coppejans et al., 2020). The origin of the giant optical flares observed from AT2022tsd in the months after the event is unclear, but timescale arguments constrain the scale of the emitting region to the outer regions of an accretion disc, for a stellar to intermediate mass black hole central engine (Ho et al., 2023a).
6 Conclusions
In this paper, we have investigated whether failed supernova in (very) massive star core-collapse events are plausible progenitors for Luminous Fast Blue Optical Transients (LFBOTs). Adopting a reasonable spread for the metallicity-dependent cosmic star formation history, we explored the rates, metallicities and expected observational characteristics of these events by identifying suitable progenitors in binary population synthesis models. Our main conclusion are as follows,
1.
The rate of core-collapse events producing black holes more massive than 30–40 M⊙ is similar to the observed LFBOT rate,
2.
The expected metallicities of the progenitors of these black holes are comparable, albeit slightly lower and formally inconsistent, with the observed metallicities of LFBOT host galaxies. However, for most LFBOTs the measured metallicity is that of the host galaxy as a whole. Even for AT2018cow, whose metallicity is most discrepant with the predicted distribution, the spatial resolution of this measurement allows for the possibility of lower metallicity pockets on smaller spatial scales,
3.
Given that the progenitors of BHs with mass greater than 30–40 M⊙ have envelopes which are H and He poor, this is consistent with the early-time spectra of LFBOTs,
4.
We show that the mass loss rates and stellar wind speeds associated with these stellar models correspond to dense circumstellar media, consistent with the lower end of LFBOT radio observations, but struggling to reproduce the most dense environments inferred. However, as bpass only runs until the end of core carbon burning, possible large mass loss phases occurring after carbon burning could explain the high-densities,
5.
The long-lived UV emission seen in AT 2018cow can be reproduced in the failed supernova interpretation given black hole masses M⊙ and several solar masses of fallback accretion.
6.
The failed supernova interpretation of LFBOTs, and models for super-kilonovae from massive collapsars, share similar progenitors and black hole masses. Therefore, if LFBOTs are massive star core-collapse events, they may constitute a non-negligible contribution to the r-process budget of galaxies.
We therefore conclude that the failed supernova scenario is a plausible explanation for at least some LFBOTs, but it hinges on the measured metallicities at the LFBOT site and circumstellar densities. The true feasibility of this progenitor channel clearly depends on the detailed physics of very massive star core-collapse, late-stage mass loss, and the amount of mass which falls back or is ejected.
Acknowledgements.
AAC acknowledges support through the European Space Agency (ESA) research fellowship programme. P.G.J. is supported by the European Union (ERC, Starstruck, 101095973, PI Jonker). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.This work made use of v2.2.1 of the Binary Population and Spectral Synthesis (bpass) models as described in Eldridge et al. (2017) and Stanway & Eldridge (2018). This work has made use of ipython (Perez & Granger, 2007), numpy (Harris et al., 2020), scipy (Virtanen et al., 2020); matplotlib (Hunter, 2007), Seaborn packages (Waskom, 2021) and astropy,(https://www.astropy.org) a community-developed core Python package for Astronomy (Astropy Collaboration et al., 2013; Price-Whelan et al., 2018).