### v2

# Luminous Fast Blue Optical Transients as very massive star core-collapse events

Abstract
Context. Luminous Fast Blue Optical Transients (LFBOTs) are rare extragalactic events of unknown origin. Tidal disruptions of white dwarfs by intermediate mass black holes, mergers of black holes and Wolf-Rayet stars, and failed supernovae are among the suggested explanations.
Aims. In this paper, we explore the viability of very massive star core-collapse events as the origin of LFBOTs. The appeal of such a model is that the formation of massive black holes via core collapse may yield observational signatures that can match the disparate lines of evidence that point towards both core-collapse and tidal disruption origins for LFBOTs.
Methods. We explore the formation rate of massive black holes in binary population synthesis models, and compare the metallicities of their progenitors with the observed metallicities of LFBOT host galaxies. We further examine the composition, mass loss rates and fallback masses of these stars, placing them in the context of LFBOT observations.
Results. The formation rate of black holes with masses greater than 30–40 M⊙ is found to be similar to the observed LFBOT rate. The stars producing these black holes are biased to low metallicity (Z0.3 Z⊙), are H and He-poor and have dense circumstellar media. However, some LFBOTs have host galaxies with higher metallicities than predicted, and they typically have denser local environments (plausibly due to late stage mass loss not captured in the models). We find that long-lived emission from an accretion disc (as implicated in the prototypical LFBOT AT 2018cow) can only be produce in these events under maximal disc mass and angular momentum conditions.
Conclusions. We conclude that (very) massive star core-collapse is a plausible explanation for at least some LFBOTs, but faces challenges. The preferred progenitors for LFBOTs in the failed supernova interpretation overlap with those predicted to produce super-kilonovae. We therefore suggest that LFBOTs are promising targets to search for super-kilonovae, and that they may contribute non-negligibly to the r-process enrichment of galaxies.
Key Words.:
Supernovae: general – Stars: black holes – Stars: massive1 Introduction
Luminous Fast Blue Optical Transients (LFBOTs), or Cow-like transients after the prototypical AT 2018cow (2018ApJ...865L...3P; 2019MNRAS.484.1031P), are rapidly evolving transients with peak optical-ultraviolet luminosities rivalling superluminous supernovae (SLSNe). The early optical-UV spectra are well described by a hot, featureless blackbody (T20000 K), lacking H and He lines initially. They have decay rates of 0.3 magnitudes/day, ruling out a significant contribution from 56Ni-powered emission. They are also accompanied by bright X-ray (and at late times, radio) emission. The late-time radio emission is attributed to self-absorbed synchrotron emission from an expanding blastwave in a dense circumstellar medium (CSM).
A modest sample of LFBOTs has now been identified in addition to AT 2018cow: ZTF18abvkwla (AT2018lug, 2020ApJ...895...49H), CSS161010 (2020ApJ...895L..23C; 2024ApJ...977..162G), ZTF20acigmel (AT2020xnd, 2021MNRAS.508.5138P; 2022ApJ...926..112B), AT2020mrf (2022ApJ...934..104Y), AT2022tsd (2023RNAAS...7..126M), AT2023fhn (2024MNRAS.527L..47C; 2024AandA...691A.329C) and AT2024wpp (2025MNRAS.537.3298P; 2025ApJ...993L...6N; 2025arXiv250900952N), as well as the slower-evolving and radio-faint, but otherwise similar, AT2024puz (2025ApJ...995..228S) and several other candidates which lack extensive multi-wavelength follow-up. Some transients display very blue colours, evolve rapidly at early times and produce bright X-ray emission, similar to LFBOTs, but are followed by broad lined type Ic SNe (e.g. the Einstein Probe-discovered transients EP2404014A and EP250108A, 2025ApJ...982L..47V; 2025ApJ...988L..14E; 2025ApJ...988L..13R). However, these events likely do not share the same progenitors as LFBOTs, given that they successfully launched supernovae, and we henceforth focus solely on the SN-less, Cow-like LFBOTs, whose origins remain unknown. The volumetric rate of LFBOTs is estimated to be 0.1% of the core-collapse SN rate (2023ApJ...949..120H).
Two categories of model to explain LFBOTs have emerged as most likely (e.g. 2019MNRAS.484.1031P). The first invokes tidal disruption events (TDEs) of H and He poor, compact stars (e.g. a white dwarfs) by intermediate mass black holes (IMBHs, 2019MNRAS.487.2505K). This model is challenged primarily due to the dense CSM observed in LFBOTs. Although such a environment is possible in TDEs involving stars disrupted by supermassive black holes, it is less clear whether white dwarf plus IMBH TDEs could produce similar circumstellar density profiles (2019ApJ...872...18M; 2024ApJ...974...67L). The second category of proposals are broadly themed around stellar mass black hole accretion, either in mergers/collisions (2022ApJ...932...84M; 2025Ap&SS.370...11G; 2025ApJ...986...84T; 2025arXiv251009745K) or core-collapse events (e.g. 2015MNRAS.451.2656K; 2019ApJ...872...18M; 2019MNRAS.485L..83Q; 2022MNRAS.513.3810G).
At least some sufficiently massive stars are expected to directly form black holes without launching a supernova. The landscape of explodability is thought to be complex, with neutron star and some black hole-forming events capable of launching successful supernovae, and other black hole-forming core-collapses resulting in direct black hole formation without any significant ejecta (e.g. 2003ApJ...591..288H; 2020ApJ...890...51E). Observational evidence for ‘quiet’ black hole formation comes from a lack of massive stars M⊙ in pre-supernovae imaging, and the disappearance of supergiant stars without evidence for a supernova (2009MNRAS.395.1409S; 2014ApJ...791..105W; 2015MNRAS.453.2885R; 2017MNRAS.468.4968A; 2017MNRAS.469.1445A; 2024ApJ...964..171B).
Some LFBOT models invoke such a scenario, where a massive star collapses to a black hole without a successful supernova (a ‘failed supernova’), followed by the accretion of some fraction of the remaining (‘leftover’) material onto the nascent BH (e.g. 2014ApJ...781..119P; 2015MNRAS.451.2656K; 2018MNRAS.476.2366F; 2019MNRAS.485L..83Q). The failed supernova hypothesis resembles early collapsar gamma-ray burst (GRB) models (e.g. 1993ApJ...405..273W; 1999ApJ...524..262M; 1999ApJ...526..152F), which also feature accreting black holes born in core-collapse, but without a successful (or a weak) supernova. In LFBOTs there is no evidence for a supernova, but unlike in collapsar GRBs, no gamma-ray emission has been observed, and the outflows are only mildly relativistic (e.g. 2020ApJ...895L..23C).
While some fast evolving optical transients can be explained solely with circumstellar interactions (2022ApJ...926..125P), in LFBOTs there is evidence for central engine activity. AT 2018cow showed short timescale variability in its light-curve (2019ApJ...871...73H; 2019ApJ...872...18M; 2022NatAs...6..249P), as well as UV and X-ray emission lasting at least several years post-explosion (2022MNRAS.512L..66S; 2023MNRAS.519.3785S; 2023MNRAS.525.4042I; 2023ApJ...955...43C; 2024ApJ...963L..24M; 2025MNRAS.544L.108I). Magnetar central engines have been proposed (2019ApJ...878...34F; 2020ApJ...888L..24M; 2022ApJ...935L..34L; 2024ApJ...963L..13L), but struggle to explain both the early-time emission and late-time optical/UV plateau (2023ApJ...955...43C). Assuming instead that LFBOTs are black hole (BH) powered, estimates for the BH mass in such a scenario have been derived from quasi period oscillations (2022NatAs...6..249P) and accretion disc modelling (2023MNRAS.525.4042I; 2024ApJ...963L..24M; 2024A&A...691A.228C; 2025MNRAS.544L.108I) yielding masses which span the upper end of the stellar mass BH distribution into the intermediate BH mass regime. The observation of late-time, short duration giant optical flares from AT 2022tsd (2023Natur.623..927H) may also be consistent with a highly variable accretion rate, either around the black hole formed in the event or a companion black hole (2024ApJ...972L..17L). Similar flaring behaviour has not been found in AT20224wpp (2025ApJ...993...76O).
Black hole mass estimates in LFBOTs are high, and they have thus far been discovered in non-nuclear regions of low-metallicity, star-forming galaxies. The aim of this paper is to explore other observational consequences - for example in terms of volumetric rates and the metallicities of their hosts - if LFBOTs are indeed powered by massive BH formation in failed supernovae. In Section 2 we compare population synthesis predictions for the rate of BH formation as a function of mass with the observed LFBOT rate. We then investigate in Section 3 whether the metallicity bias associated with the formation of the most massive stellar mass BHs is consistent with LFBOT environments. We explore in Section 4 whether the selected stellar models can plausibly reproduce the key characteristics of LFBOT emission, with a discussion and conclusions following in Sections 5 and 6.
2 Black hole formation and LFBOT event rates
We first ask the question: what is the volumetric formation rate of stellar-mass, core-collapse black holes with mass MBH,min, and which value of MBH,min best reproduces the LFBOT rate? To explore this, we couple population synthesis with a model for the metallicity-dependent cosmic star formation rate history.
For the population synthesis we use the Binary Population and Spectral Synthesis (bpass) models (v2.2.1, 2017PASA...34...58E; 2018MNRAS.479...75S). These are a publicly available grid of pre-calculated binary stellar evolution models, in which the primary evolution is followed in detail, while the secondary is evolved with the rapid evolution prescriptions of 2002MNRAS.329..897H. Models are provided at 13 metallicities and at each metallicity, each stellar model is weighted according to observed binary parameter distributions (2017ApJS..230...15M). We adopt the model set with a 2001MNRAS.322..231K broken power-law initial mass function (IMF), with a minimum mass of 0.1 M⊙ and a maximum of 300 M⊙, a break mass of 0.5 M⊙ and a slope of () below (above) the break. At each metallicity , the weighting of each model corresponds to the number of systems resembling that model expected in a stellar population of M⊙. For full details about the binary stellar models and the code, we refer the reader to the bpass v2 release paper 2017PASA...34...58E.
For the metallicity-dependent cosmic star formation rate history, CSFH, we adopt the model of 2006ApJ...638L..63L. We seed the population synthesis models according to this CSFH such that the volumetric birth rate of each system (bpass model) at each redshift and metallicity is proportional to the product of the star formation rate density SFRD and the model weighting (defined by the IMF and binary parameter distributions as described above).
We adopt the standard bpass assumptions that core-collapse occurs when, at the end of the model, the total mass exceeds 1.5 M⊙, the CO core mass exceeds 1.38 M⊙ and the ONe core mass is non-zero. Remnant masses are a pre-calculated output of the models. They are determined by injecting 1051 erg of energy, calculating the mass in the outer layers which can be lifted to infinity by this energy injection, and taking the remaining mass as the remnant mass (for full details see 2004MNRAS.353...87E; 2017PASA...34...58E). We note that this energy injection is likely an overestimate for BH-forming events and means we have a large amount of mass potentially available for fallback accretion (this is addressed further in the discussion). We classify all core-collapse remnants with M3 M⊙ as neutron stars, and those heavier as black holes. We assume that all NS-forming events produce successful supernovae, and that BH-forming events do not. Pair-instability supernovae (PISNe) are deemed to occur if the final carbon-oxygen core mass exceeds 60 M⊙ and the final helium core mass is 133 M⊙. In this case no remnant is left behind (2002ApJ...567..532H). Pulsational pair instability supernovae (PPISNe) are accounted for by adopting the remnant mass prescription of 2019ApJ...887...53F for final carbon-oxygen masses between 38 and 60 M⊙, following 2023MNRAS.520.5724B. The true landscape of explodability and remnant formation is more complex, and our assumptions neglect islands of explodability above the mass regime where neutron stars are formed, but capture the broad behaviour predicted by many works.(e.g. 2009MNRAS.395.1409S; 2009ARA&A..47...63S; 2012ApJ...749...91F; 2020ApJ...890...51E; 2020MNRAS.492.2578S; 2020MNRAS.499.2803P; 2021ApJ...909..169K; 2022MNRAS.511..903P; 2025A&A...695A..71L; 2025A&A...695A.122U; 2025A&A...700A..20M). These assumptions also naturally reproduce observed (binary) remnant mass distributions (e.g. 2014ApJ...785...28K; 2023MNRAS.520.5724B; 2023A&A...676A..31D; 2025arXiv250818083T). Nevertheless, we acknowledge the possible impacts of these simplifying assumptions in Section 5.3.
The remnant formation time for each stellar model is the seed time, plus the age of the model at core-collapse (although for these short-lived stars, this is negligible compared with cosmological timescales). In this way, we construct a metallicity-weighted supernova rate and BH formation rate as a function of redshift. Since all (confirmed) LFBOTs thus far have redshifts , we compare the mean rate of supernovae and BH-forming events over this redshift range, for different values of MBH,min. Specifically, we calculate the mean ratio of the supernova rate and the formation rate of BHs with mass MBH,min or greater. As the form of CSFH is uncertain, we repeat the process with two extreme variations of the CSFH. These are defined by shifting the 2006ApJ...638L..63L model by 0.2 dex in 12log(O/H), which captures the approximate envelope of possible CSFHs found by 2019MNRAS.488.5300C.
The results are shown in Figure 1. Using the fiducial CSFH we find that the birth rate of BHs at with masses 38 M⊙ is 0.1% of the CCSN rate, matching LFBOTs. The progenitor temperatures, luminosities and radii pre-collapse are provided in Figure 7 of Appendix A. For a CSFH of the same form but peaking at a metallicity 0.2 dex lower in 12log(O/H), we find a higher BH mass threshold of 41 M⊙ is required, because there are more high-mass BHs being formed overall with respect to CCSNe. The reverse is true for a CSFH which peaks 0.2 dex higher, yielding a threshold of 33 M⊙. Therefore, varying the CSFH changes the absolute number of BHs being born (although not the general shape of the BH mass distribution, 2023ApJ...948..105V), and the number formed with respect to CCSNe.
Above mass cuts of 45 M⊙, the ratio of birth rates remains constant, because the cut has moved into the PISN mass gap, and only black holes with masses above the gap are contributing. In summary, we find that the birth rate of BHs heavier than several tens of Solar masses is similar to the LFBOT rate, and that this result is robust against uncertainties in the CSFH.
3 Metallicity dependence and LFBOT host galaxies
$a$$a$footnotetext: IFU metallicity measurement, resolves the host giving a local (few hundred parsec scale) value.$b$$b$footnotetext: Metallicity derived from the host’s total stellar mass using equation 3 of 2004ApJ...613..898T. Uncertainties are solely propagated from uncertainties on the galaxy stellar mass, and do not include scatter in the relation.
| LFBOT | Z/Z⊙ | Reference |
| AT 2018cow | 0.470.01 | 2020MNRAS.495..992La |
| ZTF18abvkwla | 0.37 | 2020ApJ...895...49H |
| CSS161010 | 0.15 | 2024ApJ...977..162G |
| ZTF20acigmel | 0.15 | 2021MNRAS.508.5138Pb |
| AT2020mrf | 0.27 | 2022ApJ...934..104Y |
| AT2023fhn | 0.080.02 | 2024AandA...691A.329C |
| AT2024wpp | 0.53 | 2025MNRAS.537.3298Pb |
Given a black hole mass cut MBH,min which yields a BH formation rate matching the observed LFBOT rate, for an assumed CSFH, we can next ask: what is the (CSFH weighted) metallicity distribution of the progenitors of these black holes, and how does this compare with the host galaxies of LFBOTs? We collate the available information in the literature on LFBOT host metallicities in Table 1. Four of the seven have host-averaged metallicities derived from spectral energy distribution (SED) fitting. In one case (AT 2018cow) there is integral field unit data available, enabling a more local (resolution of a few hundred parsec, 2020MNRAS.495..992L) measurement of the metallicity. For the remaining two, no direct measurement has been published, so we obtain an approximate value through the galaxy mass-metallicity relation (2004ApJ...613..898T). We adopt Z by mass fraction for Solar metallicity, but note that this specific choice has no impact on the comparisons or analysis since both the observed and synthetic metallicities are scaled by the same factor.
In Figure 2 we compare the number of black holes more massive than MBH,min born per M⊙ of star formation, weighted by the mean metallicity distribution for each of the three CSFHs adopted, with the measured distribution of LFBOT host or environmental metallicities. There is a strong preference for BHs with masses greater than 38 M⊙ to be born in stellar populations with metallicity less than 0.3 Z⊙. Below this threshold, the distributions are relatively constant: this reflects that the fact that (i) we are tailoring our mass cut to yield a BH formation rate that is 0.1% of the CCSN rate for each CSFH, and (ii) that the CCSN rate itself is not substantially varying with the CSFH variations. This is because NS formation efficiency, unlike BH formation efficiency, does not sharply decline at higher metallicities (e.g. 2025ApJ...979..209V).
It is immediately apparent that LFBOT environments sample the ZZ⊙ region. The formation of BHs with MMBH,min across this range instead samples metallicities ZZ⊙. A KS-test between the LFBOT host values and the fiducial M⊙ distribution returns , formally rejecting the null hypothesis that the observed host values are consistent with the synthetic distribution. However, we note that the host metallicities are far from uniformly determined - with methods ranging from resolved IFU spectroscopy, probing the local environment within the host, to host-integrated metallicites to the mass-metallicity relation. This likely introduces errors/scatter which has not been accounted for, and host-integrated measurements will naturally be higher than the lowest metallicity environments within each host. There may also be pockets of low metallicity gas within and around star-forming galaxies, driving low-Z star-formation (e.g. 2015A&A...582A..78M; 2023MNRAS.520..879M). Nevertheless, the current information suggesting that LFBOTs can occur in environments with metallicities above the strong cut-off predicted is a challenge to the model. This is most notable for AT2018cow, whose metallicity measurement is integrated across the immediate (few hundred pc) environment (2020MNRAS.495..992L), but even in this case variations on smaller scales are possible.
4 Expectations for the emission
4.1 H and He-poor spectra
LFBOTs display H and He poor spectra at early times (e.g. 2018ApJ...865L...3P). The progenitor star, or star being tidally disrupted, therefore must have a He or He-poor composition. In the TDE scenario, a white dwarf is favoured (2019MNRAS.487.2505K). For a BH-stellar merger, the stellar object must be a Wolf-Rayet star (2022ApJ...932...84M). In the scenario explored in this paper, the median H and He mass fractions in the ‘leftover mass’222I.e. the stellar mass leftover after remnant formation. In a successful supernova with no fallback, this would be the ejecta mass. of the models selected as progenitors of BHs with M⊙ are X0.003 and Y0.08 respectively (where the uncertainties are defined by the 16th and 84th percentiles). The distributions of Xsurf and Ysurf are shown in Figure 3. 69% of our selected models are classified as Wolf-Rayet stars given the Xsurf values and high surface temperatures of log10(T/K)4.45 (2017PASA...34...58E). The outer layers of these stars therefore meet the criteria for being H and He poor, and successfully exploding stars with such low X and Y surface mass fractions are expected to produce type Ic supernovae (i.e. SNe without detectable H and He lines, e.g. 2012MNRAS.424.2139D; 2013MNRAS.436..774E; 2020MNRAS.491.3479C).
4.2 Dense circumstellar media and radio emission
A key characteristic of LFBOTs is their slow-rising, luminous radio emission, attributed to self-absorbed synchrotron emission from a blast-wave expanding through the circumstellar medium (CSM). These blast-waves are mildly relativistic (e.g. 2020ApJ...895L..23C), and probe the CSM at radii of 1016–1017 cm. LFBOT radio emission is well described by blast-wave propagation through dense CSM profiles defined by values of 1–100, in units of M⊙ yr-1/1000 km s-1. For Wolf-Rayet wind speeds of order 1000 km s-1, radii of 1016–1017 cm corresponds to mass loss in the final decades before explosion. We calculate the mean final values for the selected models, noting that bpass only runs until the end of core carbon burning. Mass loss rates are provided as a bpass output. We adopt standard bpass wind speed prescriptions, namely those of 2001A&A...369..574V, and 2000A&A...360..227N for Wolf-Rayets (we refer the reader to 2017PASA...34...58E; 2022MNRAS.515.2591C, for more details). We find a median final = (0.04) M⊙ yr-1/1000 km s-1. The full distribution of is shown in Figure 5. The bulk of our progenitor mass-loss rate and wind speed values lie below the values inferred from LFBOT radio observations, which span the range 1-100 in these units (e.g. 2022ApJ...932..116H; 2022ApJ...926..112B, under the assumption of and ). We note that mass loss in the final decades before core-collapse may be significant and eruptive, plausibly explaining the non- profiles observed (e.g. 2014ARA&A..52..487S). However, because the stellar evolution models used end after core carbon burning, a full modelling of the CSM around these stars on the relevant scales is beyond the scope of this paper.
4.3 Long-lived ultraviolet plateaus
Finally, we ask whether the selected stellar models can produce a long-lived tail of UV emission lasting at least several years after the event, as observed in AT2108cow (2022MNRAS.512L..66S; 2023MNRAS.519.3785S; 2023ApJ...955...43C; 2023MNRAS.525.4042I; 2025MNRAS.544L.108I). This has been interpreted as due to a long-lived accretion disc, reminiscent of the emission seen in TDEs (2022ApJ...932...84M; 2023MNRAS.525.4042I; 2025MNRAS.541..429M). To see whether our models can produce AT 2018cow-like UV plateaus through this mechanism, we first examine the remaining mass in the selected bpass models which survives outside the nascent BHs (i.e., the pre-core-collapse mass minus the initial BH mass). In Figure 4 we show 2D histograms of these masses at each metallicity, for the fiducial 2006ApJ...638L..63L CSFH. For more details on the BH mass distribution in bpass, we refer the reader to 2017PASA...34...58E, 2022MNRAS.511.1201G and 2023MNRAS.520.5724B. More detail on the progenitors, in terms of temperature, luminosity and radius at the end of the models (i.e. the end of core carbon burning) is provided in Section A of the Appendix.
As outlined in Section 2, bpass calculates BH masses by injecting a standard supernova energy of 1051 erg to determine the mass of the material which is unbound. The remainder is the remnant mass. In a failed supernova, the ‘leftover’ material is not successfully ejected. It may therefore be reasonable to assume that all of the leftover material falls back and is accreted instead, in line with other prescriptions for final BH masses in the massive to very massive star regime, which have close to a one-to-one relation between the pre-core-collapse mass and final remnant mass (after fallback) for stripped envelope progenitors (2012ApJ...749...91F; 2023MNRAS.520.5724B). With the BH mass MBH, leftover mass Mleftover, total pre-collapse stellar mass M⋆, radius R⋆ and rotational velocity v⋆ as inputs, we determine an initial disc radius as follows,
where and are the fraction of the pre-collapse angular momentum and leftover mass which go into the disc, respectively. A derivation of Equation 1 is provided in Appendix B. Since stellar rotation is not explicitly tracked in bpass, we have assumed the pre-collapse rotational velocity to be 10% of the breakup (critical) velocity (massive stars in low metallicity environments such as the Magellanic Clouds are observed to be rotating at 10–20% of their critical velocity e.g. 2013A&A...560A..29R; 2015A&A...580A..92R). We proceed to use the model of 2020MNRAS.492.5655M; 2025MNRAS.544.2225M to calculate UV light-curves for every selected stellar model. The final parameter required (beyond the disc mass, black hole mass and radial scale) is the initial evolutionary timescale of the disc. We follow standard SS73 theory and take , where is a dimensionless factor, and is the disc aspect ratio. Anticipating a thick disc (for this likely initially super-Eddington flow) we sample in the range , inducing scatter in our simulated light curves. This fixes all parameters in the disc model.
We adopt two example scenarios, and show the predicted UV lightcurves in Figure 6 for the F225W Wide Field Camera 3 (WFC3) Hubble Space Telescope filter. In first two, only 10% of the available angular momentum and mass is transferred to the disc. In the second extreme case, all of the angular momentum and mass goes into the disc. We acknowledge that the details details of supernova fallback, such as its asphericity and the fraction which falls back versus being ejected, are not captured in the models or our simple parametrisation. This motivated our choice of two scenarios which nevertheless represent a wide range of possible fallback masses and angular momenta, including the hypothetical maximum case. The total leftover mass is also uncertain (and depends on the injected energy at core-collapse), for a discussion of the impact of this we refer the reader to Section 5.3.
The resulting light-curves can only match the observed late-time F225W flux for AT 2018cow if a large fraction of the leftover mass and angular momentum are transferred to the disc (see also Omand et al. in prep). The initial disc radii predicted from these models are considerably smaller than the 30-40 R⊙ inferred from the late-time fits to AT2018cow (2023MNRAS.525.4042I; 2024ApJ...963L..24M). This is because the majority of the progenitors are compact stars (e.g. Wolf-Rayets, see Figure 7 of Appendix A). However, the discs spread rapidly (as t2/3), reaching 10 R⊙ within a day and 100 R⊙ within 1000 days. We note that the factors facc and fj,disc cancel, such that the two cases adopted return the same initial disc radii. The growth of the outer disc radius will be less pronounced if angular momentum is transported away from the disc by winds (potentially likely at early times), see Appendix B. We find that covers a wide range of 0.5-106 days. Only for the shortest of these timescales can the same accretion mechanism plausibly explain the rapid early-time rise in LFBOT lightcurves. Otherwise, we acknowledge that some other process must drive the initial rise and peak.
The failed SN interpretation of LFBOTs is predicated on the notion that some non-negligible fraction of the total stellar mass does not participate in the prompt BH formation and instead falls back onto the nascent BH. If the core-collapse of such massive stars actually results in the entire stellar mass being promptly engulfed by the event horizon, there will no material to accrete, ruling out failed supernovae as a viable channel for LFBOTs. However, in our fiducial models this is not the case (see Figure 4, as well as a discussion of the explodability assumptions in Section 5.3).
5 Discussion
5.1 LFBOT environments
In the failed SN scenario, we would expect LFBOTs to be strongly associated with star-forming regions, as the short-lived massive progenitors are not expected to travel far from their birth sites. Although the LFBOT sample is small, it appears as though LFBOTs occupy diverse environments within their hosts, with at least two events at moderate to large host-normalised offsets (2024MNRAS.527L..47C; 2025MNRAS.537.3298P). The apparent preference for LFBOTs to occur in low metallicity environments may help in this regard: in general, galactic outskirts have lower metallicity than their inner regions (e.g. 2012A&A...540A..56P; 2025ApJ...978L..39J). However, a similar argument could be made for e.g. core-collapse gamma-ray bursts, which are typically not observed at large offsets. Regardless of offsets, the host metallicities of LFBOTs extend above the sharp cut-off predicted for very massive star core-collapse events (see Sec 3). Furthermore, metallicities from SED fitting should be treated with caution (e.g. 2025A&A...695A..86N). For example, the mass-metallicity relation for the host of AT 2023fhn suggests a metallicity of 0.5 Z⊙, versus the fitted value of 0.1 Z⊙. In any case, the metallicity distributions as presented in Fig. 2 are inconsistent, but the possibility of low-Z pockets on spatial scales smaller than probed by current measurements has the potential to resolve this discrepancy.
5.2 Initial mass function
We have adopted model weightings with an initial mass function (IMF) slope above 0.5 M⊙ of -2.35. Two other high-end slopes are available with the bpass outputs: -2.00 and -2.70. If the -2.70 slope is adopted, fewer high-mass black holes are born, and the cut-off MBH,min drops to 30 M⊙ such that the BH formation rate remains consistent with the LFBOT rate (0.1% of the CCSN rate). The number of neutron star progenitors also decreases, such that the SN rate decreases, reducing the amount by which MBH,min has to decrease in order to maintain 0.1% of the CCSN rate. If -2.00 is adopted, there are more black holes and the cut-off moves up to 40 M⊙, around the lower edge of the PISN mass gap. This range (i.e. 30–40 M⊙) represents the uncertainty in the minimum black hole mass required to match the LFBOT rate which arises from the choice of IMF, and is slightly larger than the uncertainty arising solely from the choice of CSFH. Nevertheless, the broad conclusion is that the formation rate of black holes with masses greater than 30–40 M⊙ is consistent with the LFBOT rate.
5.3 Explodability and remnant mass assumptions
The results in this paper depend on the BH masses formed and the mass available in the ‘leftover’ material for fallback. The models adopted calculated BH masses by injecting 1051 erg of energy and determining the mass which which remains bound. If the actual energy injection in failed events is 1051 erg (as expected, e.g. 2018MNRAS.476.2366F; 2023MNRAS.526..152K) then the BH masses will be greater. However, the ejecta/leftover masses are typically not more than 10% of the BH mass, so the most they could increase by is 10, which would not significantly change our conclusions. Ultimately, the models predict similar final remnant masses to other widely used prescriptions in the high mass regime (e.g. 2012ApJ...749...91F, see also Figure 8 of Appendix A).
There is also the question of which stars explode and which undergo a failed collapse. Our assumption that all BH-forming events fail is an upper limit on the rate of failed SNe. The predicted rate of all failed SNe under this assumption is 0.2 of the CCSN (NS) formation rate, which is compatible with upper limits on the rate of failed SNe from searches for faint, long-lived red transients (thought to be a likely outcome for failed SNe in the absence of any additional energy injection from a disc, 2022MNRAS.514.1188B).
However, we are likely overestimating the failed SN rate, given recent hydrodynamical simulations and suggestions that there may not be a missing red supergiant problem (e.g. 2025ApJ...987..164B; 2025ApJ...979..117B). If failed SNe only occur in the 12-15 M⊙ (ZAMS mass) window where failed events are reliably predicted to occur in these simulation (and assuming NS formation everywhere outside this range unless a PISN occurs), then we would expect a ratio of failed events to CCSNe of 0.15. The lower masses (higher up the IMF) counteract the narrower mass range, producing a similar ratio to our fiducial 30–40 M⊙ BH mass cut. Such a scenario could therefore reproduce the LFBOT rate but not the inferred BH masses. The prospect of incorporating more physically motivated prescriptions for remnant and ejecta mass distributions in the framework of population synthesis (e.g. 2020MNRAS.499.2803P; 2025A&A...700A..20M) could help refine results such as these in future.
5.4 A potential source of r-process nucleosynthesis
Massive black holes and their accretion discs, such as those discussed here, have also been predicted as possible sites of ‘super-kilonovae’. In this scenario, the density in a high accretion disc becomes high enough to trigger r-process nucleosynthesis (2019Natur.569..241S; 2022ApJ...941..100S; 2024PhRvD.110h3024D; 2024ApJ...968...14R; 2025arXiv250315729A). The possibility of such a channel has implications for where and when r-process elements are produced in the Universe (2025ApJ...982..144N). Given the black hole and accreted masses involved in the failed supernova interpretation of LFBOTs, they may be promising sites for r-process production. The volumetric rate of LFBOTs in the redshift range considered in this work (2023ApJ...949..120H, 0.1% of the CCSN rate, or 70 yr-1 Gpc-3) is compatible with the rate of super-kilonovae predicted by 2022ApJ...941..100S of 10–100 yr-1 Gpc-3. We have shown that the failed-supernova interpretation of LFBOTs naturally favours events forming black holes greater than M⊙, while super-KNe have been predicted in events producing black holes greater than M⊙ (2022ApJ...941..100S). If the true rate of LFBOTs is just a factor of few lower than the current 70 yr-1 Gpc-3 best estimate, then the BH mass cut required to match the LFBOT rate moves up to the pair instability mass gap (see Figure 1), and we have a 1:1 mapping between LFBOT and super-KN progenitors. In any case, there is surely substantial overlap in the populations. Furthermore, a near-infrared excess - a hallmark of radiative reprocessing by r-process elements - was observed on a timescale of weeks following both AT 2018cow (2019MNRAS.484.1031P) and AT2024wpp (2025MNRAS.537.3298P). This has been described either as a dust echo or due to free-free opacity effects outside the optical photosphere in both events (2023ApJ...944...74M; 2025ApJ...993L...6N), but no spectroscopic characterisation of this feature in an LFBOT has yet been obtained. Future (James Webb Space Telescope) near-infrared spectroscopic observations could potentially distinguish between an r-process element-forming super-kilonova and a featureless dust echo or free-free emission. However, in order for a UV-emitting disc and IR-emitting super-kilonova to be simultaneously observable, an anisotropic geometry which allows simultaneous viewing of both photospheres would be required.
5.5 Further implications
There are several other observations associated with LFBOTs which any model invoked to explain them must satisfy, assuming that LFBOTs are homogenous class. For example, extremely high velocity (), blue-shifted hydrogen lines were observed in the CSS161010 (2024ApJ...977..162G), perhaps most easily explained as out-flowing streams of material in a tidal disruption event, although these have only been observed in CSS161010 thus far. Polarimetric observations have found that LFBOTs start highly polarised and trend towards low polarisation on a timescale of days (2023MNRAS.521.3323M; 2025MNRAS.537.3298P). This can plausibly be explained in a failed supernova context by (i) an accretion disc producing the high early-time polarisation and (ii) a rapid transition to an spherical outflow (2025MNRAS.537.3298P), perhaps caused by outflows from highly super-Eddington accretion (e.g. 2020ApJ...895L..23C). The origin of the giant optical flares observed from AT2022tsd in the months after the event is unclear, but timescale arguments constrain the scale of the emitting region to the outer regions of an accretion disc, for a stellar to intermediate mass black hole central engine (2023Natur.623..927H).
6 Conclusions
In this paper, we have investigated whether failed supernova in (very) massive star core-collapse events are plausible progenitors for Luminous Fast Blue Optical Transients (LFBOTs). Adopting a reasonable spread for the metallicity-dependent cosmic star formation history, we explored the rates, metallicities and expected observational characteristics of these events by identifying suitable progenitors in binary population synthesis models. Our main conclusion are as follows,
1.
The rate of core-collapse events producing black holes more massive than 30–40 M⊙ is similar to the observed LFBOT rate,
2.
The expected metallicities of the progenitors of these black holes are comparable, albeit slightly lower and formally inconsistent, with the observed metallicities of LFBOT host galaxies. However, for most LFBOTs the measured metallicity is that of the host galaxy as a whole. Even for AT2018cow, whose metallicity is most discrepant with the predicted distribution, the spatial resolution of this measurement allows for the possibility of lower metallicity pockets on smaller spatial scales,
3.
Given that the progenitors of BHs with mass greater than 30–40 M⊙ have envelopes which are H and He poor, this is consistent with the early-time spectra of LFBOTs,
4.
We show that the mass loss rates and stellar wind speeds associated with these stellar models correspond to dense circumstellar media, but which nevertheless only overlap with the lower end of LFBOT radio observations, and struggle to reproduce the typical circumstellar densities inferred. However, as bpass only runs until the end of core carbon burning, possible large mass loss phases occurring after carbon burning could explain the high-densities,
5.
The long-lived UV emission seen in AT 2018cow can be reproduced in the failed supernova interpretation given black hole masses M⊙, but this requires (i) several solar masses of fallback accretion and (ii) a high fraction of the pre-collapse stellar angular momentum to be transferred to the disc,
6.
The failed supernova interpretation of LFBOTs, and models for super-kilonovae from massive collapsars, share similar progenitors and black hole masses. Therefore, if LFBOTs are massive star core-collapse events, they may constitute a non-negligible contribution to the r-process budget of galaxies.
We therefore conclude that the failed supernova scenario is a plausible explanation for at least some LFBOTs, but it hinges on the measured metallicities at the LFBOT site and the origin of the dense circumstellar media. The true feasibility of this progenitor channel clearly depends on the detailed physics of very massive star core-collapse, late-stage mass loss, and the amount of mass which falls back or is ejected.
Acknowledgements.
We thank the referee Brian Metzger for their insightful comments on this manuscript. AAC acknowledges support through the European Space Agency (ESA) research fellowship programme. P.G.J. is supported by the European Union (ERC, Starstruck, 101095973, PI Jonker). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.This work made use of v2.2.1 of the Binary Population and Spectral Synthesis (bpass) models as described in 2017PASA...34...58E and 2018MNRAS.479...75S. This work has made use of ipython (2007CSE.....9c..21P), numpy (2020Natur.585..357H), scipy (2020NatMe..17..261V); matplotlib (2007CSE.....9...90H), Seaborn packages (Waskom2021) and astropy,(https://www.astropy.org) a community-developed core Python package for Astronomy (astropy:2013; astropy:2018).
Appendix A Additional figures
In this Appendix we provide two additional figures. Figure 7 shows the progenitor temperatures, luminosities and radii for the fiducial selection of BH progenitors where the BH mass exceeds 38 M⊙. Figure 8 shows the carbon-oxygen core versus BH masses for the selected models.
Appendix B Disc radii
To calculate the initial disc radii - for input to the model of 2020MNRAS.492.5655M, 2025MNRAS.541..429M - we first define M⋆=MMleftover. The initial disc radius can be estimated from angular momentum and mass conservation as the star is “turned” (from some process we do not specify) into a Keplerian flow. We therefore need an estimate of the maximum angular momentum available to the disc. On energetics grounds we can write for the pre-collapse star,
where is the maximum pre-collapse stellar rotational velocity. The corresponding specific angular momentum of the star is approximately . Substituting into equation B.1 (and ignoring constants of order unity) we find
and therefore
The angular momentum in the accretion disc has to be less than or equal to the total pre-collapse angular momentum, therefore
Where and is the initial disc radius. We therefore find an absolute maximum333In principle the initial disc radius could be larger than this if the disc material somehow got all of the initial stellar angular momentum but not all of the leftover mass. We ignore this possibility here. initial disc radius of
in the extreme case that (i) the star is maximally rotating and (ii) the total pre-collapse angular momentum is transferred entirely to the disc. Since and we can introduce factors to allow for sub-critical rotation velocities , and the fraction of angular momentum which goes into the disc , as follows:
Finally, we adopt for the pre-collapse rotational velocity and introduce the fraction of leftover mass which is accreted , yielding
as adopted in Section 4.3.
The disc then must expand, as angular momentum is conserved while the mass of the flow drops (as some material is added to the black hole). To be precise, the quantity
is conserved. The mass in the disc drops with time in a self similar fashion, as the accretion rate
at large times (e.g. Pringle81), such that
and thus
Of course, this expansion of the disc can be suppressed if angular momentum is lost from the disc – i.e., if angular momentum is transported to large radii by winds. It is important therefore that the radii at large times (in this model) are larger than that inferred from observations.